OpenStudy (anonymous):

how do you do this problem...: find z x w z=6(cos3π/2+i3π/2) w=12(cos5π/6+isin5π/6)

OpenStudy (anonymous):

do you know?

OpenStudy (anonymous):

ultiply rthem i know that!

OpenStudy (tkhunny):

Do you have De Moivre?

OpenStudy (anonymous):

yeah the formula?

OpenStudy (tkhunny):

Right. It should show you the way. \((a\cdot cis(\theta))\cdot (b\cdot cis(\phi)) = ??\)

OpenStudy (anonymous):

ok thanks